OpenStudy (anonymous):

let f be the function defined by the following. f(x)= sinx, x<0 x^2, 0 (lessthanorequalto) x< 1 2-x, 1(lessthanorequalto)x<2 x-3 x(greaterthanorequalto)2 For what value of x is f NOT continious? 0 only 1 only 2 only 0 and 2 only or 0.1.2

myininaya (myininaya):

\[\lim_{x \rightarrow 2^-}f(x)=\lim_{x \rightarrow 2^-}(2-x)=0\] \[\lim_{x \rightarrow 2^+}f(x)=\lim_{x \rightarrow 2^+}(x-3)=-1\] left does not equal right limit not continuous at 2

myininaya (myininaya):

do the same thing for 0 and 1 see if left limit=right limit

OpenStudy (anonymous):

what do u mean do the same for 0 and 1?

myininaya (myininaya):

check the left and right limit at 0 and 1 like i did above for 2

OpenStudy (anonymous):

(2-x) okay so for limit as x approaches 0, it would be 2..and for 1, the answer is one

myininaya (myininaya):

left limit at 0 is sin(0)=0 right limit at 0 is (0)^2=0 since left limit=right limit and f(0)=0 then the function is continuous at 0